Find the value of -za/2 that corresponds to a level of confidence of 91.28 percent. 1) -1.71 2) 1.71 3) 0.0436 4) -1.36Question 12 (6 points) Saved Find the z-score corresponding to the given value and use the z-score to determine whether the v
Answer #1

0 given lind of lind of confidenil 91.84 1-25 91.28 0:9128 loo 22 -0.9128 2= 0.0%72 7 010437 be Now the value of - 2212 can f7- scare is defined as 7 = X-M Here x = 15-osilond H17.5 stungs and ta 2:1 seconds - Z-score = 15.0-17:5 2:1 12- -sane =-1.2